# Robot Navigation with Entity-Based Collision Avoidance using Deep Reinforcement Learning

### v1

@@ 3.3 Algorithm Parallel Deep V-learning
The value network is trained by the temporal-difference method with standard experience replay and fixed target network techniques. However, we improve it by using parallelism. We run several agents in several environments in parallel which helps us to collect the experience much faster than in the previous work. We call this approach “Parallel Deep V-learning", the pseudocode is shown in Algorithm 1 and Algorithm 2.
The approach that we present in Algorithms 1 and 2 is a bit similar to asynchronous methods for deep reinforcement learning Mnih et al. (2016). However, the key difference here is that we do not do asynchronous updates of neural network. Instead, we accumulate state and value pairs into Experience replay in parallel and then we make a centralized update using the whole replay buffer with random sampling. In our setup, the episode duration is much longer than the gradient learning step, so the bottleneck of the original method was in running episodes. So using parallel deep V-learning solves the problem even without asynchronous model updates. In the future, we can also test our algorithm with asynchronous model updates.
@@
1:Run imitation learning with demonstration $D$
2:Update value network $V$ with $D$
3:Initialize target value network $\hat{V}\leftarrow V$
4:Initialize experience replay memory $E\leftarrow D$
5:Initialize parallel processes number $N$
6:while episode $<$do
7:     Initialize a multiprocessing Pool of size $N$
8:     for $k$ = episode, episode + $N$ do in parallel:
9:         Run the function $RUN\_EPISODE(k)$
10:     end for
11:     Sample random minibatch tuples from $E$
12:     Update value network $V$ by gradient descent
13:     if episode $\mod$ UpdateInterval = 0 then
14:         Update target value network $\widetilde{V}\leftarrow V$
15:     end if
16:     Update $episode\leftarrow episode+N$
17:end while
18:return $V$
@@
1:function RUN_EPISODE
2:     Initialize environment
3:     Initialize robot with current model $V$
4:     Initialize random sequence $s_{0}^{j}$
5:     Initialize temporary buffer $S$
6:     repeat
7:         $\mathbf{a}_{t}\leftarrow\begin{cases}\text{RandomAction}(),\qquad\qquad\epsilon\\ \arg\max_{\mathbf{a}_{t}\in\mathbf{A}}R(\mathbf{s}_{t}^{j},\mathbf{a}_{t})+\\ \qquad+\gamma^{\Delta t\cdot v_{\text{pref}}}V(\mathbf{s}_{t+\Delta t}^{j}),\quad 1-\epsilon\end{cases}$
8:         Make step using $\mathbf{a}_{t}$
9:         Store tuple $(s_{t}^{j},a_{t},r_{t},s_{t+\Delta t}^{j})$ in $S$
10:         Update time $t\leftarrow t+\Delta t$
11:     until $s_{t}$ is a terminal state or $t\geq t_{\max}$
12:     if Need to update memory then
13:         for $t=0$, length($S$do
14:              Get $r_{t}$, $s_{t}^{j}$ and $s_{t+1}^{j}$ from $S$
15:              Set $value_{t}=r_{t}+\gamma^{\Delta t\cdot v_{\text{pref}}}\hat{V}(s_{t+1}^{j})$
16:              Store ($s_{t}$, $value_{t}$) in $E$
17:         end for
18:     end if
19:end function
@@ 3.4 Implementation Details
We assume holonomic kinematics for the robot, i.e., it can move in any direction. The action space consists of 80 discrete actions: 5 speeds exponentially spaced between (0, $v_{pref}$] and 16 headings evenly spaced between $[0,2\pi)$.
The hidden units of functions $\phi(\cdot)$, $\psi_{h}(\cdot)$, $\psi_{\alpha}(\cdot)$, $f_{v}(\cdot)$ are (300, 200), (200, 100), (200, 200), (300, 200, 200) respectively.
We implemented the policy in PyTorch and trained it with a batch size of 100 (for each episode) using stochastic gradient descent. For imitation learning, we collected 3000 episodes of demonstration using ORCA and trained the policy 50 epochs with a learning rate of 0.01. For reinforcement learning, the learning rate is 0.001 and the discount factor $\gamma$ is 0.9. The exploration rate of the $\epsilon$-greedy policy decays linearly from 0.5 to 0.1 in the first 5000 episodes and stays at 0.1 for the remaining 5000 episodes.
@@ 4.1 Simulator
Figure 1: Simulator example.
We train and test our models using our simulator. The example of one step of an episode is shown in Figure 1. The robot, adults, bicyclists, children, and static obstacles are shown. The goals of each entity (excluding static obstacles) and the robot’s goal are marked as stars. Each dynamic entity and its goal is enumerated. We use a square crossing scenario, where all the agents are randomly positioned on a square and their goal positions are chosen randomly on the opposite side of the square. Also, we use a circle crossing scenario, where all the agents are randomly positioned on a circle and their goal positions are chosen on the opposite side of the circle.
@@ 4.2 Simulation Setup
The simulated agents: adults, children, and bicyclists are controlled by Optimal Reciprocal Collision Avoidance (ORCA) van den Berg et al. (2011). The sizes and speed of agents are generated from Uniform distribution to introduce diversity. The parameters of the distribution depend on the entity type. The distribution parameters are chosen so that, on average, adults are slightly larger and faster than children, and cyclists are slightly larger and faster than adults. Additionally, we account for the presence of slow cyclists and fast children, as well as for the presence of small cyclists and large children. Consequently, it is not possible to accurately determine the type of object solely based on its size and speed.
We use a square crossing scenario, where all the agents are randomly positioned on a square of size 11m $\times$ 11m. For simplicity, in our experiments, we represent static objects as several stationary pedestrians. However, in future work, we plan to incorporate static grid processing into our method. We compare the results with the state-of-the-art method SARL Chen et al. (2019) and its modification SARL-GP which is the SARL method with a reward function that uses a Goal Proximity component.
We evaluate the effectiveness of the proposed model in the invisible setting. In this setting, the robot is invisible to other agents. This setting could be sometimes more complicated than the setting when other agents react to robot movement, as it tests the “worst case" scenario when the robot should avoid the collision itself. We also think this is a good setting for comparing methods since robots are often overlooked by humans, cyclists, and vehicles due to their size. As a result, the simulated agents react only to other agents but not to the robot.
Models are trained for 3000 episodes using imitation learning and for 50000 episodes using Parallel Deep V-learning algorithm 1. We validate models every 1024 episodes on a Validation dataset of size 500. After train step for each method, we select the model that achieved the highest reward metric on the validation dataset. We conduct a final comparison using a Test dataset consisting of 1000 episodes.
Figure 2: Collision Rate, Success Rate, Reward and Weighted Score metrics on Train and Validation data.
Figure 3: Density histograms of danger distances on Test dataset.
@@ 4.3 Results
The results of the train and validation steps are shown in Figure 2. Firstly, based on the Success Rate metric we can see that the SARL method completely fails to learn policy that can reach the goal. The SARL model can only learn a policy that successfully avoids collisions. SARL has the lowest Collision Rate metric. As our main task is to provide an algorithm that achieves the goal (while avoiding obstacles), we do not consider the SARL model in further analysis, since it fails to do such a task completely.
@@
We can see that both methods SARL-GP and EB-CADRL (Ours) successfully converge during 50000 train episodes and get to a plateau approximately from 30000 episodes. We can see that EB-CADRL outperforms other methods both on train and validation using the Collision Rate metric and Weighted Score metric. We compute the Weighted Score metric using the following formula:
$Weighted\ Score$ $=SR-CR(A)-4.0\cdot CR(C)$ (18)
$\quad-2.0\cdot CR(B)-0.5\cdot CR(O)$
@@
The reward metric in this case cannot be used to compare different models as each model has its own reward function. However, it indicates how the method converges and could be used to get the episode for early stopping based on results on the validation dataset.
In Table 1 we provide the result of testing each model on test data which has a size of 1000 episodes. For comparison, we use the following metrics: SR - success rate, CR - collision rate, CR(A) - collision rate with adults, CR(B) - collision rate with bicycles, CR(C) - collision rate with children, CR(O) - collision rate with static obstacles, Time - average time to reach the goal, DD(A) - mean danger distance with adults, DD(B) - mean danger distance with bicycles, DD(C) - mean danger distance with children, DD(O) - mean danger distance with static obstacles. We assume that danger starts when the distance between the robot and the entity is less than 30cm.
Method SR CR CR(A) CR(B) CR(C) CR(O) Time DD(A) DD(B) DD(C)
SARL 0 0.009 0.003 0.004 0.002 0 NaN 0.175 0.173 0.185
SARL-GP 0.681 0.207 0.027 0.062 0.029 0.089 29.32 0.13 0.113 0.131
EB-CADRL 0.681 0.068 0.02 0.018 0.002 0.028 29.45 0.118 0.154 0.165
@@
Table 1: Metrics on Test data. We indicate best metric value in comparison between SARL-GP and EB-CADRL (since SARL fails to learn to navigate to the goal), differce less than 0.5% is not indicated.
SARL has a success rate of 0, meaning it fails to reach the goal in all attempts. Both SARL-GP and EB-CADRL have a success rate of 0.681, indicating they are equally successful in reaching the goal. SARL has the lowest collision rate, but it’s important to note that its success rate is zero, indicating it did not navigate successfully enough to encounter many obstacles. EB-CADRL has a significantly lower collision rate compared to SARL-GP, indicating better overall safety. EB-CADRL has lower collision rates with bicycles and static obstacles compared to SARL-GP and is on par with SARL in terms of collisions with children. It also has a lower collision rate with adults compared to SARL-GP, although slightly higher than SARL. SARL-GP and EB-CADRL have similar average times to reach the goal, indicating that EB-CADRL’s improved safety does not come at the expense of significantly increased travel time. The time to reach the goal for SARL is invalid since SARL has a 0 Success Rate, so it is indicated as NaN in the table. EB-CADRL maintains a safer distance from adults compared to both SARL and SARL-GP, but it is closer to bicycles and children compared to SARL-GP. However, it still achieves a balance between maintaining safe distances and minimizing collisions. To sum up, EB-CADRL is the better method overall because it achieves a good balance between reaching the goal successfully, minimizing collisions, and maintaining safe distances. It demonstrates the best safety performance while maintaining efficiency, making it suitable for robot navigation in environments with dynamic and static entities of different types.
In Figure 3 we provide the histograms of the distribution of danger distances of SARL-GP and EB-CADRL methods. We can see that the EB-CADRL model tends to be safer with bicyclists and children.
@@ 4.4 Ablation Experiments
To evaluate the impact of including the agent type as a feature, we conducted an ablation study by training and testing two models: one incorporating the agent type and the other excluding it. While the model can infer some aspects of the agent type through features such as radius and velocity, these features alone do not provide a clear distinction between different agent types. For instance, a small, slow-moving cyclist or a large, fast-moving human could mislead the model. Therefore, we hypothesize that explicitly adding the agent type to the feature set could enhance the model’s performance. Both models in this experiment utilize a new reward function and are based on the EB-CADRL method.
For ablation experiments, we use circle and square crossing scenarios. The radius of the circle is 6m and the size of the square is 13m $\times$ 13m.
To test models during ablation experiments we use the following metrics. Reward is the reward function that we formulated in 3.1. Success measures the percentage of times the robot successfully reached its goal without collisions. Collision metrics (Adult, Bicycle, Child, Obstacle) measure the frequency of collisions with different entities. Danger time is the mean time during which the robot was in a potentially dangerous state (close to collision). A lower value means the robot spent less time in dangerous zones in which there is a risk of getting into a collision. Apart from that, the long presence of the robot in such areas may cause discomfort to surrounding pedestrians and other entities which is not desirable. Time is the average time it took for the robot to reach its goal. A lower time indicates that the robot was able to complete its task more quickly.
No entity type With entity type
Reward 0.037 0.052
Success 0.448 0.501
Collision (Bicycle) 0.004 0.002
Collision (Child) 0.001 0.001
Collision (Human) 0.015 0.016
Collision (Obstacle) 0.002 0.005
Danger time 0.913 0.746
Time 39.078 32.857
Table 2: Metrics on Test data during ablation experiments.
The results of the ablation experiment on the Test dataset are shown in Table 2. The reinforcement learning model with entity type outperforms the one without entity type in several key aspects. The model with entity type achieved a higher reward and a better success rate, indicating superior overall performance. It also registered fewer collisions with bicycles, a shorter time to reach the goal, and less time in danger. However, it is worth noting that the model without entity type had fewer collisions with adults and obstacles. Overall, the inclusion of entity type in the model appears to enhance its performance in most aspects.
Figure 4: Reward and success rate metrics on Train and Validation data during ablation experiments.
5 Discussion
While our EB-CADRL method shows promising results, several limitations remain. The simulation environment may not fully capture real-world complexities such as varying terrains and unpredictable human behaviors. The predefined entity types and simplified reward function might limit the model’s adaptability to more nuanced scenarios. Moreover, the current implementation assumes holonomic kinematics, which may not be applicable to all types of robots. Future work should focus on real-world testing, incorporating advanced perception systems, developing adaptive reward functions, optimizing computational efficiency, extending the model to non-holonomic robots, and exploring multi-robot coordination. Addressing these areas will enhance the robustness, safety, and efficiency of autonomous robot navigation in dynamic environments.

### v2

@@ 4.3 Algorithm Parallel Deep V-learning
The value network is trained by the temporal-difference method with standard experience replay and fixed target network techniques [15]. However, we improve it by using parallelism. We run several agents in several environments in parallel which helps us to collect the experience much faster than in the previous work. We call this approach “Parallel Deep V-learning”, the pseudocode is shown in Algorithm 1 and Algorithm 2.
On a 24-core CPU, our proposed Parallel Deep V-learning algorithm was 12 times faster than conventional Deep V-learning [4, 6] in both training and evaluation. This significant speed-up allows for rapidly training more sophisticated models in environments with numerous dynamic agents and static obstacles.
@@
1:Run imitation learning with demonstration $D$, update value network $V$ with $D$
2:Initialize: target value network $\hat{V}\leftarrow V$, experience replay memory $E\leftarrow D$, parallel processes number NN
3:while episode << M do
4: Initialize a multiprocessing Pool of size NN
5:for $k$ = episode, episode + NN do in parallel:
6:  Run the function RUN_EPISODE(k)RUN\_EPISODE(k)
7:end for
8: Sample random minibatch tuples from $E$
9: Update value network $V$ by gradient descent
10:if episode $\mod$ UpdateInterval = 0 then
11:  Update target value network $\widetilde{V}\leftarrow V$
12:end if
13: Update $episode\leftarrow episode+N$
14:end while
15:return $V$
@@
1:function RUN_EPISODE(k)
2: Initialize: environment, robot with current model $V$, random sequence $s_{0}^{j}$
3: Initialize temporary buffer $S$
4:repeat
5:  $\mathbf{a}_{t}\leftarrow\begin{cases}\text{RandomAction}(),\qquad\qquad\qquad\qquad\qquad\qquad\epsilon\\ \arg\max_{\mathbf{a}_{t}\in\mathbf{A}}R(\mathbf{s}_{t}^{j},\mathbf{a}_{t})+\gamma^{\Delta t\cdot v_{\text{\text{pref}}}}V(\mathbf{s}_{t+\Delta t}^{j}),\quad 1-\epsilon\end{cases}$
6:  Make step using $\mathbf{a}_{t}$
7:  Store tuple $(s_{t}^{j},a_{t},r_{t},s_{t+\Delta t}^{j})$ in $S$
8:  Update time $t\leftarrow t+\Delta t$
9:until $s_{t}$ is a terminal state or $t\geq t_{\max}$
10:if Need to update memory then
11:  for $t=0$, length($S$) do
12:   Get $r_{t}$, $s_{t}^{j}$ and $s_{t+1}^{j}$ from $S$
13:   Set $value_{t}=r_{t}+\gamma^{\Delta t\cdot v_{\text{\text{pref}}}}\hat{V}(s_{t+1}^{j})$
14:   Store ($s_{t}$, $value_{t}$) in $E$
15:  end for
16:end if
17:end function
@@ 4.4 Implementation Details
We assume holonomic kinematics for the robot, i.e., it can move in any direction. The action space consists of 81 discrete actions: a stop action (0 velocity) plus combinations of 5 speeds exponentially spaced between (0, $v_{\text{pref}}$] and 16 headings evenly spaced between $0,2\pi)$.
The neural network architecture consists of four multi-layer perceptrons (MLP). An MLP’s structure is denoted by the tuple $(d_{\text{in}},d_{1},\dots,d_{\text{out}})$, which specifies the number of neurons in the input, hidden, and output layers, respectively. The dimensions of these MLP components are $\phi_{g}$: (17, 300, 200), $\psi_{h}$: (200, 200, 100), $\psi_{\alpha}$: (400, 200, 200, 1), and $f_{v}$: (106, 300, 200, 200, 1). The input dimension 17 to $\phi_{g}$ represents the joint state of the robot’s state (6), agent’s state (7), and entity type information (4), while the 106 input dimensions to $f_{v}$ combine the robot’s state (6) with the attention-weighted feature representation of all agents (100).
We implemented the policy in the deep learning framework PyTorch [16] and trained it with a batch size of 100 (for each episode) using stochastic gradient descent. For imitation learning, we collected 3000 episodes of demonstration using ORCA and trained the policy 50 epochs with a learning rate of 0.01. For reinforcement learning, the learning rate is 0.001 and the discount factor $\gamma$ is 0.9. The exploration rate of the $\epsilon$-greedy policy decays linearly from 0.5 to 0.05 in the first 25000 episodes and stays at 0.05 for the remaining episodes.
All input features, including the robot’s state, entity states, and entity types, would be provided by localization and perception systems on a real robot. For our experiments, this information is obtained directly from the simulator’s state, while the implementation of these systems on a physical robot is assumed to rely on existing algorithms [17] and is outside the scope of this research.
@@ 5 Experiments
5.1 Reward setup
In this subsection we specify the reward function setup for our experiments, where the robot navigates in the environment containing 4 different types of entities: adults, bicycles, children, and static obstacles. The number of each entity type varies, with experiments involving between 3 and 9 of each, creating crowded scenarios with over 26 agents and obstacles.
We denote the entity types as follows: Adult ($A$), Bicycle ($B$), Child ($C$), and Obstacle ($O$). The penalties for collisions are defined as:
[…]
The rationale behind the collision penalties is based on the predictability of the entity’s behavior and the severity of the consequences of a collision. Static obstacles are the most predictable and have the least severe consequences, resulting in the smallest penalty (-0.5). Adults exhibit predictable behavior with moderate collision consequences, resulting in a penalty (-1.0). Cyclists, due to their higher speed and the potential for severe consequences such as falling off a bicycle, incur a higher penalty (-1.5). Children are the most unpredictable and vulnerable, leading to the most severe consequences in the event of a collision resulting in the highest penalty (-2.0).
Let $d_{\text{disc}}(e)$ be the discomfort distance (in meters) for entity $e$:
[…]
Let $p_{\text{disc}}(e)$ be the discomfort penalty factor for entity $e$:
[…]
The discomfort distance $d_{\text{disc}}(O)$ and discomfort penalty factor $p_{\text{disc}}(O)$ are omitted, since the robot does not cause discomfort for static obstacles.
@@ 5.2 Simulator
We train and test our models using our simulator. The example of one step of an episode is shown in Figure 1. The robot, adults, bicyclists, children, and static obstacles are shown. The goals for the robot and dynamic entities are marked as stars. Each entity and the corresponding goals are enumerated.
Figure 1: A simulator snapshot at $\text{Time}=10.0$s from the start, showing 8 adults, 8 bicyclists, 8 children, and 3 static obstacles. Static obstacles may overlap.
We use a square crossing scenario, where all the agents are randomly positioned on a square and their goal positions are chosen randomly on the opposite side of the square. Also, we use a circle crossing scenario, where all the agents are randomly positioned on a circle and their goal positions are chosen on the opposite side of the circle.
@@ 5.3 Simulation Setup
Simulated agents (adults, children, bicyclists) are controlled by Optimal Reciprocal Collision Avoidance (ORCA) [2]. To introduce diversity, each agent’s size and speed are drawn from a uniform distribution specific to its type. The distributions are designed to overlap, for instance, some children may be faster than some adults, making it impossible to determine an agent’s type from its size and speed alone. The experiments are conducted in an 11m $\times$ 11m square crossing scenario. Each episode contains one robot, 8 adults, 8 cyclists, 8 children, and 3 static obstacles, all randomly positioned. For simplicity, static obstacles are modeled as stationary pedestrians, with plans to incorporate static grid processing in future work. We compare five navigation methods: CADRL [6], LSTM-RL [8], the state-of-the-art SARL [4], a modification called SARL-GP (which adds Time (6) and Goal Proximity (7) rewards to SARL), and our EB-CADRL. All evaluations use an “invisible robot” setting, where other agents do not perceive or react to the robot. This tests the robot’s collision avoidance capabilities in a worst-case scenario that is also realistic, as small robots are often overlooked by people.
Figure 2: Collision rate, success rate, reward, weighted score metrics on the training and validation data. The validation step is performed every 1024 episodes.
Models are trained for 3000 episodes with imitation learning, followed by 50000 episodes using the Parallel Deep V-learning algorithm 1. During training, we validate the models every 1024 episodes on a 500-episode validation set. For the final comparison, we select the best-performing model for each method based on its validation performance and evaluate it on a 1000-episode test set.
@@ 5.4 Results
Figure 2 shows the training and validation results. SARL fails to learn a goal-reaching policy, learning only to avoid collisions, and is thus excluded from further analysis. CADRL and LSTM-RL methods also have poor performance. Additionally, CADRL exhibits an unacceptably high collision rate toward the end of training.
@@
Both SARL-GP and EB-CADRL (ours) successfully converge within 50000 training episodes, reaching a plateau at about 30000 episodes. EB-CADRL outperforms the other methods in both training and validation, as measured by collision rate and weighted score metrics. We compute the Weighted Score $(WS)$ metric using the following formula:
[…]
The coefficients in the Weighted Score metric can be adjusted according to business needs and safety priorities. The current coefficients are chosen to penalize more heavily collisions with more vulnerable entities (children and bicyclists).
@@
The reward metric cannot be used to compare different models as each model has its own reward function. However, it indicates how the method converges and could be used to get the episode for early stopping based on results on the validation dataset.
@@
Table 1: Metrics on test data. We indicate best metric value in comparison between CADRL, LSTM-RL, SARL-GP and EB-CADRL
Metric SARL CADRL LSTM-RL SARL-GP EB-CADRL
WS -0.05 -0.36 -0.66 0.37 0.60
SR 0 0.05 0.01 0.68 0.68
CR 0.017 0.175 0.268 0.207 0.068
CR(A) 0.002 0.027 0.0485 0.027 0.020
CR(B) 0.008 0.099 0.125 0.062 0.018
CR(C) 0.007 0.047 0.093 0.029 0.002
CR(O) 0 0.003 0.002 0.089 0.028
Time NaN 28.62 27.46 29.32 29.45
DD(A) 0.177 0.108 0.083 0.130 0.118
DD(B) 0.174 0.101 0.077 0.113 0.154
DD(C) 0.170 0.113 0.083 0.131 0.165
In Table 1 we provide the result of testing each model on test data which has a size of 1000 episodes. For comparison, we use the following metrics: SR - success rate; CR - overall collision rate; CR(A), CR(B), CR(C), CR(O) - collision rates with adults, bicycles, children and static obstacles, respectively; Time - average time to reach the goal; DD(A), DD(B), DD(C) - mean danger distances with adults, bicycles and children, respectively. We assume that danger starts when the distance between the robot and the entity is less than 30cm.
SARL has a success rate of 0, indicating that it failed to reach the goal in all attempts. CADRL and LSTM-RL have very low success rates, making these methods nearly as ineffective as SARL. SARL-GP and EB-CADRL have a success rate of 0.68, indicating they are similarly effective in reaching the goal. SARL has the lowest collision rate but failed to reach any goals, indicating it did not navigate successfully enough to encounter many agents and obstacles. EB-CADRL has a significantly lower collision rate compared to CADRL, LSTM-RL and SARL-GP, indicating better safety. Also, EB-CADRL has lower collision rates with adults, bicycles and children compared to CADRL, LSTM-RL and SARL-GP.
Figure 3: Density histograms of danger distances on the test dataset
SARL-GP and EB-CADRL have similar average times to reach the goal, indicating that EB-CADRL’s improved safety does not come at the expense of increased travel time. Time metric for SARL is indicated as NaN, since SARL has a 0 success rate. Although the LSTM-RL shows the shortest average time to reach the goal, this result is less meaningful due to the method’s low success rate.
EB-CADRL maintains safer distances from bicycles and children compared to CADRL, LSTM-RL and SARL-GP. Moreover, it significantly outperforms all other methods based on the weighted score metric.
Figure 3 shows the density histograms of danger distances for SARL-GP and EB-CADRL. EB-CADRL provides enhanced safety when interacting with bicyclists and children, resulting in significantly fewer cases where the robot operates within 0–0.2 meters of these agents compared to SARL-GP.
To sum up, EB-CADRL is the superior method overall because it achieves a good balance between reaching the goal successfully, minimizing collisions, and maintaining safe distances, which makes it efficient and safe for navigation in environments with dynamic and static entities of different types.
@@ 5.5 Ablation Experiments
Table 2: Metrics on test data during ablation experiments
SR CR(A) CR(B) CR(C) CR(O) Reward Time Danger time
No entity type 0.448 0.015 0.004 0.001 0.002 0.037 39.078 0.913
With entity type 0.501 0.016 0.002 0.001 0.005 0.052 32.857 0.746
Figure 4: Reward and success rate metrics on the training and validation datasets during ablation experiments.
To evaluate the impact of including the agent type as a feature, we conducted an ablation study by training and testing two models: one that incorporates the agent type and the other that excludes it. Although the model can infer some aspects of the agent type through features such as radius and velocity, these features alone do not provide a clear distinction between different agent types. For instance, a small, slow-moving cyclist or a large, fast-moving adult could mislead the model. Therefore, we hypothesize that explicitly adding the agent type to the feature set will enhance the model’s performance. Both models in this experiment utilize a new reward function and are based on the EB-CADRL.
The results of the ablation experiment on the test dataset are shown in Table 2. In addition to the metrics defined in 5.4, we compute the Danger time metric: the mean time (in seconds) the robot spent in a potentially dangerous state. A lower value indicates safer navigation and less discomfort caused to other entities.
The model incorporating the entity type outperforms the one excluding it in several key aspects. It achieved a higher reward and better success rate, indicating superior overall performance. Additionally, it registered fewer bicycle collisions, reached the goal faster, and spent less time in danger. However, the model excluding the entity type had slightly fewer collisions with adults and obstacles. The reward and success rate metrics on the training and validation data are shown in Figure 4.